# CSF: Fixed-outline Floorplanning Based on the Conjugate Subgradient Algorithm Assisted by Q-Learning

### v1

@@ III-D The Constraint Graph
It is worth mentioning that a properly constructed CG pair (HCG, VCG) can be directly used to translate into a floorplan where all modules are mutually disjoint.
IV The CSA-based Fixed-outline Floorplanning Algorithm
As presented in Section III-A, the stages of global floorplanning and legalization can be implemented by addressing (2) and (3), respectively. Due to the non-smooth characteristics, we propose a CSA-based fixed-outline floorplanning algorithm (CSF). In the CSF, the non-smooth optimization problems (2) and (3) are addressed by CSA [22]. Meanwhile, we also improve the performance of CSA by adaptively regulating the scaling factor of step size via Q-learning.
IV-A The Framework of CSF
The framework of CSF is illustrated by Algorithm 2. At the beginning, a solution population $Pop$ is initialized, where the orientations of modules are randomly selected and the latin hypercube sampling (LHS) is employed to generate the initial position of the module $fp_{j}$ for each individual $ind_{i}$ in the solution population [22], $p$ and $N$ represent the population size and the total number of modules respectively. The characteristic of LHS lies in that it can make the generated samples distribute as evenly as possible in the parameter space, thus covering a broader solution space and being more likely to find the global optimal solution. Then, the stages of global floorplanning and legalization are successively implemented. If a legal solution $\mathbf{FP^{*}}$ is obtained after the legalization stage, record this solution as the final solution and terminate the iteration. Otherwise, some randomly selected modules are rotated and the floorplanning process can be repeated again until the iteration budget is exhausted.
Algorithm 2 The framework of $CSF$
@@ IV-A The Framework of CSF
2:Legal floorplan $\mathbf{ind}^{*}=(\mathbf{fp}_{1},\dots,\mathbf{fp}_{N})$;
3:Generate $\mathbf{ind}_{i}=(\mathbf{fp}_{1},\dots,\mathbf{fp}_{N})(i=1,\dots,p)$ by LHS;
@@
6:     Implement global floorplanning by addressing (2);
7:     Implement legalization by addressing (3);
8:     if the solution $\mathbf{FP}^{*}$ obtained after legalization is legal then
9:         Set $\mathbf{ind}^{*}=\mathbf{FP}^{*}$;
10:         break;
11:     end if
12:     Rotate selected modules of $\mathbf{ind}_{i}(i=1,\dots,p)$;
13:     t++;
14:end while
IV-B The Conjugate Subgradient Algorithm Assisted by Q-learning
To address the fixed-outline floorplanning problem efficiently, we propose a conjugate subgradient algorithm assisted by Q-learning (CSAQ), which employs a population-based iteration scheme with the principle of CSA. Thanks to the gradient-like definition and the stochastic characteristics of subgradient, the CSA incorporates both the fast convergence of gradient-based algorithm and the global convergence of metaheuristics. The step size controlled by $c$ plays a crucial role in achieving the balance between exploration and exploitation.
Because appropriate behaviors make it easier for individuals to find the optimal strategy to obtain the maximum long-term return [28], we employ a dynamical update strategy of $c$ using a Q-learning-assisted strategy [29].
IV-B1 The Q-learning strategy
For a population of size $p$, a $p\times m$ Q-table is constructed as Table I, where each state corresponds to an individual, and $m$ actions represent the update strategies of $c$ [29]. Then, the scaling factor $c$ is updated for state $s_{j}$ according to the probability distribution
$P_{i}^{j}=\frac{Q^{k}(s_{j},a_{i})}{\sum_{l=1}^{m}Q^{k}(s_{j},a_{l})},i=1,\dots,m,$ (11)
where $Q^{k}(s_{j},a_{i})$ represents the Q-value of state-action pair $(s_{j},a_{i})$ at the $k$-th iteration. At each iteration, the Q-values are updated by (9) and the reward function $R(s_{j},a_{i})$ is confirmed by .
$R(s_{j},a_{i})=(f_{pre}-f_{post})/100,$ (12)
TABLE I: The Q-table for the update of $c$.
@@ IV-B1 The Q-learning strategy
$s_{p}$/$\mathbf{ind}_{p}$ $Q(s_{p},a_{1})$ $Q(s_{p},a_{2})$ $Q(s_{p},a_{m})$
IV-B2 The framework of CSAQ
@@
Algorithm 3 $CSAQ$
1:The objective $f$, the population $\mathbf{Pop}=\{\mathbf{ind}_{j},j=1,\dots,p\}$, the iteration gap $k_{t}$, the initial parameter $c$;
@@
3:Initialize the Q-table (Table I) uniformly;
@@
7:             Set $R(s_{j},a_{i})=0$ for all $i\in\{1,\dots,m\}$;
8:             $\mathbf{d}_{j,0}=\mathbf{0}$, $\mathbf{ind}^{0}_{j}=\mathbf{ind}_{j}$, $\mathbf{g}_{j,0}\in\partial f(\mathbf{ind}_{j}^{0})$;
@@
11:             Update $Q^{k}(s_{j},a_{i})$ according to (9);
12:             Update $c_{j}$ according to (11);
@@
14:         Calculate the subgradient $\mathbf{g}_{j,k}\in\partial f(\mathbf{ind}_{j}^{k-1})$;
@@
21:Set $\mathbf{Pop}=\{\mathbf{ind}^{k_{max}}_{j},j=1,\dots,p\}$.
IV-C The Global Floorplanning Stage
The global floorplanning stage is implemented by addressing problem (2) using the CSAQ described by Algorithm 3. The motivation is to get properly distributed modules by minimizing the penalized objective with increasing $\lambda$. The iteration process ceases if the total overlapping area of all individuals in the solution population is less than $A/v$, where $A$ is the sum of the areas of all modules and $v$ is a preset constant. The process of global floorplanning is shown in Algorithm 4.
Algorithm 4 Global Floorplanning by the $CSAQ$
@@ IV-C The Global Floorplanning Stage
3:Initialize parameters $\alpha$, $\lambda$ and $\mu$ of the objective $f_{g}$;
@@
5:    Perform Algorithm 3
6:    for $j=1$ to $p$ do
7:         $\mathbf{ind}_{j}=\mathbf{ind}_{j}^{k_{max}}$;
8:    end for
9:    Set $\lambda=q*\lambda$;
10:end while
11:$\mathbf{Pop}^{*}=\{\mathbf{ind}_{1},\dots,\mathbf{ind}_{p}\}$.
IV-D The Legalization Stage
After the global floorplanning stage, a legalization procedure is implemented to get a legal floorplan. In this paper, we propose two legalization strategies, the improved legalization based on the constraint graphs and the legalization based on the CSAQ.
IV-D1 Legalization based on the constraint graphs
The legalization based on the constraint graph starts with the construction of the HCG and the VCG. Then, the coordinates of the modules are renewed according to the constructed constraint graphs. If all modules are placed in the constrained region, we get the final floorplan. Otherwise, additional strategies are employed to get a legal floorplan. The process of eliminating out-of-bound aims to adjust a constraint relationship in the critical path of a CG. Since the process to eliminate horizontal bound excess is similar to that in the vertical direction, we only present the details of horizontal legalization.
Li et al. [12] proposed a CG-based legalization for large-scale floorplanning, where a critical relationship $A\to B$ is removed from the HCG, and a corresponding relationship is added to the VCG if it is needed. According to Li’s legalization algorithm, the modification of HCG is implemented as follows.
@@ IV-D1 Legalization based on the constraint graphs
    If $A\to B$ is compressible, it is removed from the HCG.
@@
    The critical relationship with the maximum weight is removed from the HCG, and a corresponding relationship is inserted to the VCG at the same time. The weight of $A\to B$ is computed by
    $weight(A\to B)=\begin{cases}S^{y}_{A}-h_{B}&\text{if }y_{A}\leq y_{B},\\ S^{y}_{B}-h_{A}&\text{otherwise},\end{cases}$ (13)
    where $S^{y}_{A}$ and $S^{y}_{B}$ are the respective vertical slacks of modules $A$ and $B$.
Although modification of the critical relationship with the maximum weight can reduce the vertical placement range to the greatest extent, it is sometimes not the best choice in terms of wirelength. Accordingly, we propose an improved legalization strategy ($ILG_{x}$) as presented in Algorithm 5.
Algorithm 5 $ILG_{x}(W,HCG,VCG)$
1:Width $W$ of the fixed outline, horizontal and vertical constraint graphs $(HCG,VCG)$;
@@
3:Calculate the width $W_{d}$ of the current floorplan and locate all the critical relationships in the $HCG$;
4:while $W_{d}>W$ do
@@
13:         if $y_{A}\leq y_{B}$ then
14:             Insert $A\uparrow B$ to the VCG;
15:         else
16:             Insert $B\uparrow A$ to the VCG;
17:         end if
@@
19:    Calculate the width $W_{d}$ and locate all critical relationships in the $HCG$;
@@
According to the weight values, the proposed $ILG_{x}$ sorts all incompressible relationships in descending order, and then selects $k$ of them with the greatest values as the candidate relationships. To select one of the $k$ candidate relationships to be removed from the $HCG$, we estimate the wirelength changes of the potential removals, and sort them in descending order of wirelength change. Then, we establish a probability distribution $\{Pw_{i},i=1,\dots,k\}$, where $Pw_{i}>Pw_{i+1}$ for $i=1,\dots,k-1$. In this way, the legalization of floorplan can lead to shorter total wirelength.
The vertical legalization $ILG_{y}$ can be carried out in similar way. By alternatively executing $ILG_{x}$ and $ILG_{y}$ until a legal floorplan is obtained or the maximum number of iterations $l_{max}$ is reached, we get the improved legalization algorithm based on constraint graphs presented in Algorithm 6.
Algorithm 6 Improved Legalization by the CGs
1:$W$, $H$, the solution $\mathbf{FP}=(\mathbf{fp}_{1},\dots,\mathbf{fp}_{N})$ to be legalized, the maximum number of iterations $l_{max}$;
2:The updated solution $\mathbf{FP}^{*}=(\mathbf{fp}_{1}^{*},\dots,\mathbf{fp}_{N}^{*})$;
3:Generate $HCG$ and $VCG$ according to $\mathbf{FP}$;
4:Update $\mathbf{FP}$ according to the $HCG$ and the $VCG$;
5:for $i=1$ to $l_{max}$ do
6:    Perform $ILG_{x}$ to update $\mathbf{FP}$;
7:    Perform $ILG_{y}$ to update $\mathbf{FP}$;
8:    if $\mathbf{FP}$ is legal then
9:         $\mathbf{FP^{*}}=\mathbf{FP}$;
10:         break;
11:    end if
12:end for
IV-D2 Legalization based on the CSAQ
Sun et al. [22] proposed to optimize (3) by CSA to get a legal floorplan. To improve the performance of CSA-based legalization, the CSA-Q is employed in this paper to address problem (3). As presented in Algorithm 7, the CSAQ used in the legalization stage is similar to that deployed in the global floorplanning stage, and the iteration process is terminated when either the maximum number of iterations is reached or a legal individual (floorplan) is obtained. Since the legalization process based on CSAQ sometimes gets an incompact floorplan, the constraint graphs are finally constructed to update the legalized floorplan to get a compact one.
Algorithm 7 Legalization by the CSAQ
1:The Objective $f_{l}$, the population $\mathbf{Pop}=\{\mathbf{ind}_{1},\dots,\mathbf{ind}_{p}\}$;
2:The updated solution $\mathbf{FP^{*}}=(\mathbf{fp}_{1}^{*},\dots,\mathbf{fp}_{N}^{*})$;
3:Initialize parameters $\lambda$ and $\mu$ of the objective $f_{l}$, set succ=0, initialize the Q-table (Table I) uniformly;

### v2

@@ 3.4 The Constraint Graph
Given a set of modules to be placed, a CG pair (HCG, VCG) can be translated into a floorplan where all modules are mutually disjoint. Then, by constructing a CG pair (HCG, VCG) for the result of global floorplanning and translating it into another floorplan of modules, the overlaps between modules can be eliminated and we get a compact floorplan for the modules to be placed.
4 The CSA-based Fixed-outline Floorplanning Algorithm
As presented in Section 3.1, the global floorplanning and the legalization can be implemented by addressing (8) and (9), respectively. Accordingly, we propose a CSA-based fixed-outline floorplanning algorithm (CSF), where the performance of CSA enhances by adaptively regulating the scaling factor $c$ of step size via Q-learning.
4.1 The Framework of CSF
The framework of CSF is presented as Algorithm 2. At the beginning, a solution population $Pop$ of size $p$ is initialized, where the orientations of modules are randomly generated and the latin hypercube sampling (LHS) is employed to generate the initial position vector $\mathbf{Ind}_{i}$ within the fixed-outline The characteristic of LHS lies in that it can make the generated samples distribute as evenly as possible in the parameter space, thus covering a broader solution space and thus increasing the likelihood of finding the global optimum. Then, the stages of global floorplanning and legalization are successively implemented. If a legal solution $\mathbf{Ind}^{*}$ is obtained after the legalization stage, record this solution as the final solution and terminate the iteration. Otherwise, some randomly selected modules are rotated for $\mathbf{Ind}_{i}$ and the floorplanning process can be repeated again until the iteration budget $t_{max}=10$ is exhausted.
Algorithm 2 The Framework of CSFCSF
@@ 4.1 The Framework of CSF
2:A floorplan $\mathbf{Ind}^{*}=(\mathbf{x}^{*},\mathbf{y}^{*})$;
3:Generate a population of floorplan $\mathbf{Pop}=\{\mathbf{Ind}_{1},\dots,\mathbf{Ind}_{p}\}$, where $\mathbf{Ind}_{i}=(\mathbf{x}^{i},\mathbf{y}^{i})(i=1,\dots,p)$ is the initial floorplan sampled within the fixed-outline by LHS;
@@
6:  $\mathbf{Pop}=GFloorplan(f_{g},\mathbf{Pop})$; /*Global floorplanning*/
7:  Implement legalization stage to get $\mathbf{Ind}^{*}$ /*Legalization*/
8:  if $\mathbf{Ind}^{*}$ is legal then
9:   $\mathbf{break}$;
10:  end if
11:  Rotate randomly selected modules for each individual $\mathbf{ind}_{i}(i=1,\dots,p)$;/*Random rotation of modules*/
12:  t++;
13:end while
4.2 The Conjugate Subgradient Algorithm Assisted by Q-learning
The global floorplanning and the legalization address problems (8) and (9) by CSAQ, which employs a population-based iteration scheme with the principle of CSA. Thanks to the gradient-like definition and the stochastic characteristics of subgradient, the CSA incorporates both the fast convergence of gradient-based algorithms and the global convergence of metaheuristics. The parameter $c$ of CSA (Algorithm 1) plays a crucial role in achieving the balance between exploration and exploitation, which is updated by a Q-learning-assisted strategy [29].
Table 1: The Q-table for the update of $c$.
@@ 4.2 The Conjugate Subgradient Algorithm Assisted by Q-learning
$s_{p}$$\mathbf{Ind}_{p}$ $Q(s_{p},a_{1})$ $Q(s_{p},a_{2})$ $Q(s_{p},a_{m})$
For a solution population $\mathbf{Pop}$ of size $p$, a $p\times m$ Q-table is constructed as Table 1, where each state represents an individual, and an action corresponds to an alternative value of $c$ [29]. In this paper, the $p$ and $m$ are set to 5. Subsequently, the scaling factor $c$ corresponding to state $s_{i}$ ($\mathbf{Ind}_{i}$) is updated via a probability distribution
$P_{i}^{j}=\frac{Q^{k}(s_{i},a_{j})}{\sum_{l=1}^{m}Q^{k}(s_{i},a_{l})},\quad j=1,\dots,m,$ (12)
where $Q^{k}(s_{i},a_{j})$ is the Q-value of state-action pair $(s_{i},a_{j})$ at the $k$-th iteration. At each iteration, the Q-values are updated by (10) where the reward function $R(s_{i},a_{j})$ is confirmed by
$R(s_{i},a_{j})=(f_{pre}-f_{post})/100,$ (13)
@@
Algorithm 3 $CSAQ(f,\mathbf{Pop})$
1:The objective $f$, the population $\mathbf{Pop}=\{\mathbf{Ind}_{i},i=1,\dots,p\}$;
@@
3:Initialize the parameters $k_{t}$ and the Q-table (Table 1) uniformly;
@@
7:     Set $R(s_{i},a_{j})=0$ for all $j\in\{1,\dots,m\}$, initialize $c_{j}$;
8:     $\mathbf{d}_{i,0}=\mathbf{0}$, $\mathbf{Ind}^{0}_{i}=\mathbf{Ind}_{i}$, $\mathbf{g}_{i,0}\in\partial f(\mathbf{Ind}_{i}^{0})$;
@@
11:     Update $Q^{k}(s_{i},a_{j})$ according to (10);
12:     Update $c_{i}$ according to the distribution (12);
@@
14:    Calculate the subgradient $\mathbf{g}_{i,k}\in\partial f(\mathbf{Ind}_{i}^{k-1})$;
15:    Calculate the parameter $\eta_{i,k}=\frac{\mathbf{g}_{i,k}^{T}(\mathbf{g}_{i,k}-\mathbf{g}_{i,k-1})}{\lVert\mathbf{g}_{i,k-1}\rVert_{2}^{2}}$;
16:    Calculate the conjugate direction $\mathbf{d}_{i,k}=-\mathbf{g}_{i,k}+\eta_{i,k}\mathbf{d}_{i,k-1}$;
17:    Calculate the step size $\hat{a}_{i,k}=\frac{c_{i}}{\lVert\mathbf{d}_{i,k}\rVert_{2}}$;
18:    Update the solution $\mathbf{Ind}_{i}^{k}=\mathbf{Ind}_{i}^{k-1}+\hat{a}_{i,k}\mathbf{d}_{i,k}$;
@@
21:Set $\mathbf{Pop}=\{\mathbf{Ind}^{k_{max}}_{i},i=1,\dots,p\}$.
4.3 The Global Floorplanning
The global floorplanning is implemented by addressing problem (8) using the CSAQ described by Algorithm 3. Because the weight setting of $f_{g}$ depends on the case to be investigated, we employ a dynamic setting of objective weight $\lambda$. As presented in Algorithm 4, the CSAQ is executed iteratively for the parameter setting $\lambda\leftarrow q\lambda$, where $q$ is set as $1.3$ in this paper. In this way, it tries to achieve a promising result striking a balance between wirelength and constraint violation. The iteration process ceases if the total overlapping area of all individuals in the solution population is less than $A/v$, where $A$ is the sum of the areas of all modules and $v$ is a preset constant.
Algorithm 4 $GFloorplan(f_{g},\mathbf{Pop})$
@@ 4.3 The Global Floorplanning
3:Initialize parameters $\alpha$, $\lambda$ and $\mu$ of the objective $f_{g}$ (Eqn. (8));
@@
5:  $\mathbf{Pop}=CSAQ(f_{g},\mathbf{Pop})$;
6:  $\lambda=q\lambda$;
7:end while
[…]
4.4 The Legalization
Li et al. [12] proposed a tailored legalization strategy based on the constraint graph (LA-CG), which generates promising results on large-scale floorplanning cases with soft modules. However, our preliminary experiment indicated that it does not work well on small cases without soft modules. Accordingly, we propose an improved legalization algorithm based on constraint graph (ILA-CG). Moreover, we also perform the legalization by solving problem (9). The CSAQ has been demonstrated to be an efficient optimizer, then we get an efficient legalization algorithm based on CSAQ (LA-CSAQ).
4.4.1 The legalization algorithm based on constraint graph
The LA-CG proposed by Li et al. [12] starts with the construction of the HCG and the VCG, which are then translated into a floorplan without overlap. If all modules are placed in the constrained region, we get a legal floorplan; otherwise, there are some modules placed outside the fixed outline, and constraint relationships in the critical path  [^2] The so-called critical path refers to a path composed of blocks that constrain each other in the same direction and are closely arranged [30]. of a CG are selected to be adjusted to eliminate the out-of-bound violation. Since the process to eliminate horizontal out-of-bound violation is similar to that in the vertical direction, we only present the details of horizontal legalization.
To eliminate the horizontal out-of-bound violation, the LA-CG proposes several definitions for modules and arcs in CG  [^3] Please refer to Ref. [12] for the rigorous definitions..
  • 1.
    For module $A$, the maximum adjustment range of the x-coordinate tolerated by the HCG is defined as the horizontal slack $S_{A}^{x}$; similarly, $S_{A}^{y}$ is the vertical slack defined as the maximum adjustment range of the y-coordinate.
  • 2.
    If $S_{A}^{x}=S_{B}^{x}=0$, the arc $A\to B$ (representing a relationship between $A$ and $B$) in HCG is termed as a horizontal critical relationship , which means the x-coordinates of modules $A$ and $B$ cannot be adjusted without overlapping with other modules; A vertical critical relationship $A\uparrow B$ in the VCG can be defined similarly. The arcs in the critical path of a HCG/VCG are a critical relationship.
  • 3.
    For a horizontal critical relationship $A\to B$, The weight is defined by
    $weight(A\to B)=\begin{cases}S^{y}_{A}-h_{B}&\text{if }y_{A}\leq y_{B},\\ S^{y}_{B}-h_{A}&\text{otherwise}.\end{cases}$ (14)
    The weight for vertical critical relationship $A\uparrow B$ can be defined similarly.
  • 4.
    Modules $A$ and $B$ have a compressible relationship in the horizontal/vertical direction if $A$ or $B$ may be moved horizontally/vertically such that the floorplan is more compact.
Then, the LA-CG modifies the HCG as follows.
@@ 4.4.1 The legalization algorithm based on constraint graph
    If an arc $A\to B$ is compressible, it is removed from the HCG.
@@
    The critical relationship $A\to B$ with the maximum weight is removed from the HCG, and a corresponding relationship is inserted to the VCG at the same time.
The modification of VCG is performed similarly. By modifying the critical relationship with the maximum weight, the LA-CG aims to reduce the vertical placement range to the greatest extent, employing a greedy strategy that sometimes fail to achieve the optimal modification in terms of wirelength minimization.
Although modification of the critical relationship with the maximum weight can reduce the vertical placement range to the greatest extent, it is a greedy strategy that sometimes does not get the best modification in terms of wirelength minimization. To remedy the deficiency of the LA-CG, we propose an improved legalization strategy in the $x$-direction (named as $ILG_{x}$) presented in Algorithm 5, and the $ILG_{y}$ for legalization in the $y$-direction is implemented similarly.
Algorithm 5 $ILG_{x}(\mathbf{W^{*}},HCG,VCG)$
1:Width $\mathbf{W^{*}}$ of the fixed outline, the CG pair $(HCG,VCG)$;
@@
3:Calculate the width $\mathbf{W_{d}}$ of the current floorplan and locate all the critical relationships in the HCGHCG;
4:while $\mathbf{W_{d}>W^{*}}$ do
@@
13:  end if
14:  if $y_{A}\leq y_{B}$ then
15:    Insert $A\uparrow B$ to the VCG;
16:  else
17:    Insert $B\uparrow A$ to the VCG;
@@
19:  Calculate the width $\mathbf{W_{d}}$ and locate all critical relationships in the HCGHCG;
@@
According to the weight values quantified by (14), the proposed $ILG_{x}$ gets $k$ candidate relationships by selecting critical relationships with top $k$ weight values. Then, $k$ candidate relationships are sorted in descending order by the potential wirelength increment, and we get a probability distribution $\{Pw_{i},i=1,\dots,k\}$, where $Pw_{i}>Pw_{i+1}$ for all $i\in\{1,\dots,k-1\}$. The distribution plays a significant role in the performance of the CG-based legalization, and we perform a preliminary study in Section 5.2.
By alternatively executing $ILG_{x}$ and $ILG_{y}$ until a legal floorplan is obtained or the maximum number of iterations $N_{max}$ is reached, we get the improved legalization algorithm based on constraint graph (ILA-CG) presented in Algorithm 6, trying to reduce the wirelength while legalization is implemented.
Algorithm 6 $ILA-CG(\mathbf{W,H,Pop})$
1:The population $\mathbf{Pop}=(\mathbf{Ind}_{1},\dots,\mathbf{Ind}_{p})$, width $\mathbf{W^{*}}$ and height $\mathbf{H^{*}}$ of the fixed outline;
2:The post-legalization solution/floorplan $\mathbf{Ind}^{*}=(\mathbf{x}^{*},\mathbf{y}^{*})$;
3:for $i=1,\dots,p$ do
4:  for $k=1$ to $N_{max}$ do
5:   Generate $HCG_{i}$ and $VCG_{i}$ according to $\mathbf{Ind}_{i}$;
6:   Update $\mathbf{Ind}_{i}$ according to the $HCG_{i}$ and the $VCG_{i}$;
7:   $(HCG^{\prime}_{i},VCG^{\prime}_{i})=ILG_{x}(\mathbf{W^{*}},HCG_{i},VCG_{i})$;/*Horizontal legalization*/
8:   $(HCG^{\prime}_{i},VCG^{\prime}_{i})=ILG_{y}(\mathbf{H^{*}},HCG_{i},VCG_{i})$;/*Vertical legalization*/
9:   Update $\mathbf{Ind}_{i}$ according to the $HCG^{\prime}_{i}$ and the $VCG^{\prime}_{i}$;
10:   Calculate the width $\mathbf{W_{d}}$ and height $\mathbf{H_{d}}$ of the current floorplan;
11:   if $\mathbf{W_{d}\leq W^{*}}$ and $\mathbf{H_{d}\leq H^{*}}$ then
12:     $\mathbf{Ind}^{*}=\mathbf{Ind}_{i}$;
13:     $\mathbf{return}$;
14:   end if
15:  end for
16:end for
4.4.2 The legalization algorithm based on CSAQ
Since the ILA-CG does not explicitly optimize the wirelength, it sometimes generates a legalized floorplan with poor wirelength metric. Then, Sun et al. [22] proposed to directly optimize (9) by CSA to get a legal floorplan. To further improve the performance of CSA-based legalization, the CSAQ is employed in this paper, and we get the legalization algorithm based on CSAQ (LA-CSAQ) presented in Algorithm 7. The CSAQ used in the legalization stage is similar to that deployed in the global floorplanning stage, and the iteration process is terminated when either the maximum number of iterations is reached or a legal individual (floorplan) is obtained. Optimization of problem (9) by CSAQ sometimes gets an incompact floorplan (Fig. 1(a)), which incurs unnecessary wirelength costs. Accordingly, the CG pair (HCG,VCG) is constructed to update the legalized floorplan to make the modules compactly arranged (Fig. 1(b)).
(a) The floorplan before compress.
(b) The floorplan after compress.
Figure 1: The floorplan compression implemented by CG.
Algorithm 7 $LA-CSAQ(\mathbf{f_{l},Pop})$
1:The population $\mathbf{Pop}=(\mathbf{Ind}_{1},\dots,\mathbf{Ind}_{p})$;
2:The post-legalization solution/floorplan $\mathbf{Ind}^{*}=(\mathbf{x}^{*},\mathbf{y}^{*})$;
3:Initialize parameters $\lambda$ and $\mu$ of the objective $f_{l}$, the parameter $k_{t}$, initialize the Q-table (Table 1) uniformly;